### v1

@@ A.1 Algorithm of Hydra
Algorithm 1 presents in detail the workflow of Hydra. [^4] We note that part of this workflow is the same as (or aligns with) the workflow of SFL, e.g., how part-1 of the model is processed. For this reason, we often refer to it as SFL+Hydra. The algorithm shows the steps of the three main entities (e.g., Aggregator, Server, and the Clients). In detail, the aggregator initiates the training by initializing the weights of the model, and sends the corresponding model parts to the server and the clients, i.e., weights $w_{p1}$ of part-1 are sent to clients, and weights $w_{p2_{a}},w_{p2_{b}}$ of part-2a and part-2b, are sent to the server. Then the server and the client can start training. The training takes place in $R$ rounds. In every round, the server ensures that it receives the batch updates from all clients. Note that the server ‘pulls’ a new batch request (which contains the activation from a client) using the function proc_order(). This function guarantees that the preferred processing order is followed, e.g., cyclic order (or cyclic-and-reverse, as in Appendix F.2) or random order.
@@
As we also described in Section 4, the server maintains $G$ versions of part-2b (heads). The mapping (or grouping) of each client’s intermediate data to the different part-2b heads is done by using the $u$ set, which contains the mapping of clients into groups. This set is defined by Algorithm 2, following the assignment policy described in Appendix A.2.
@@
Finally, at the end of each round, the entities aggregate the versions/replicas of the model parts. In detail, the clients send the updated weights of part-1 to the aggregator, which will compute the weights for the next round, using an aggregation function, like FedAvg. We note that Hydra’s design provides the flexibility to apply any aggregation function, as it does not alter the processing workflow. Concurrently, the server aggregates locally the replicas of the heads, e.g., the part-2b for all groups.
A.2 Group Definition in Hydra
@@
In this section, we describe in detail how clients’ data are mapped to different versions of part-2b (heads). In its core, this mapping groups clients according to the similarity in their data. We use the terms of “versions of part-2b”, “heads”, and “groups” interchangeably. Let $g\in\mathcal{G}$ represent a group, where $\mathcal{G}$ is the set of all groups, and let $G=|\mathcal{G}|$ denote the total number of groups (each group corresponds to a version of part-2b, as Figure 4 shows). Hydra’s design naturally bounds $G$ by the number of labels, i.e., $G\leq L$. This is in line with the multi-head architectures in CL, where, typically, the number of heads is equal to the number of tasks. In what follows, for simplicity, we describe the mapping/grouping of Hydra assuming that $G=L$. However, this description can be easily extended for the case where $G<L$, as we do in our performance analysis in Section 5.
@@
The assignment of clients into groups considers two factors, (factor-i) the data distribution of the clients and, specifically, the representation of the different labels in the clients’ data, and (factor-ii) repartition of clients among the groups, i.e., ensuring that all versions of part-2b receive a uniform number of updates (or, in other words, the groups have a balanced number of assigned clients). Regarding (factor-i), the goal is to group clients with similar data distributions. In this way, we avoid generating large gradients in the higher layers during iterations through clients. Recall that large gradients impose larger weight changes, which lead to overwriting previous updates, i.e., forgetting the previous knowledge. As shown in ramasesh2020anatomy (and verified in our observations), the higher layers are responsible for the large gradients that are back-propagated across the whole model. On the contrary, maintaining homogeneity within a group ensures a smooth transition between clients, minimizing significant weight changes and effectively addressing the primary challenge of CF.
Algorithm 2 Client assignment to groups
Input: num of groups $G$, set of clients $\mathcal{C}$
@@
Based on the discussion above, we propose the following approach for grouping the clients into groups and mapping them in different head layers (i.e., part-2b as shown in Figure 4). Let $\rho_{cl}$ be the number of samples of label $l$ in client $c$, where $l\in\mathcal{L}$ and $c\in\mathcal{C}$, and $\rho_{c}=\{\rho_{cl}|l\in\mathcal{L}\},\forall c\in\mathcal{C}$ be the corresponding vector for each client $c$. Furthermore, we define the binary decision variables $\mathbf{u}=\{u_{cg}\mid\forall c\in\mathcal{C},\forall g\in\mathcal{G}\}$, where $u_{cg}=1$ when client $c$ is assigned to group $g$, and $u_{cg}=0$ otherwise. By design, each client must be assigned to exactly one group, i.e.,
@@ Appendix B Variants of Split (Federated) Learning
There is a plethora of proposed definitions for SL, e.g., SplitNN vepakomma2018split or parallel SL jeon2020privacy, that mostly differ in the way the training at the server takes place and the presence of an aggregator. Among them, SplitFed thapa2022splitfed (also called Split Federated Learning [^5] We clarify that, for simplicity, the abbreviation “SFL” always refers to SplitFedv2 throughout this manuscript, unless otherwise specified.) is an emerging method for training deep Neural Network models in a system comprised of multiple IoT and mobile devices. We introduce here the most widely used variants of SplitFed and discuss their differences. Figure 7 provides an overview of each variant of SplitFed, as well as the multi-head FL paradigm chen2023flexibility. We note that the training steps (forward/backward propagation) for the three variants of SplitFed (i.e., versions 1 to 3) are similar to the ones depicted in Figure 1. Moreover, in all three variants, part-1 is handled by the clients and part-2 is handled by the server.
@@
SplitFedV1 was the first SplitFed variant proposed in thapa2022splitfed, also often referred as Parallel SL tirana2024workflow. As shown in Figure 7, there is a different version of the model parts for each client. This enables the distinct segments of the model to be trained entirely in parallel, replicating a paradigm similar to the one of FL. Additionally, the model parts are aggregated at the end of each round, and the global model is the result of concatenating the two model parts. The main drawback of this setup is the computing and memory demands on the server side, i.e., the server preserves the state of part-2 for every participating client. Moreover, the fact that the model convergence guarantees are the same as in FL, and not as in conventional SL, i.e., SplitNN vepakomma2018split, can be considered another drawback. This is because often, as we have seen in the literature, SL paradigms can outperform FL (which is also shown in Figure 16 in the case of IID data distribution).
@@
SplitFedV2 (SFL) is the second variant of SplitFed introduced in thapa2022splitfed. This version aligns more with the original (sequential) Split Learning (SL) framework, i.e., SplitNN vepakomma2018split, when compared to SplitFedV1. Unlike SplitFedV1, which operates with fully parallel model updates, SFL incorporates a shared model segment, specifically part-2, among the clients. Consequently, clients perform parallel updates on part-1, while part-2 is updated sequentially. The advantages of this approach include: (i) enhanced model accuracy relative to SplitFedV1 (FL), as shown in thapa2022splitfed and (ii) reduced server resource requirements compared to SplitFedV1. However, as discussed in this paper, SFL is susceptible to catastrophic forgetting (CF), a phenomenon that is further pronounced under heterogeneous data distributions among the clients.
@@
SplitFedV3 was introduced  gawali2021comparison and madaan2022vulnerability, and was specifically designed to address CF madaan2022vulnerability in Split Federated Learning (SFL). The training process closely mirrors that of SplitFedV1, where each client maintains its version of part-1 and part-2. However, at the end of each round, only part-2 is aggregated. Consequently, the final model features multiple input layers, corresponding to the layers that comprise part-1. Nevertheless, the process for selecting a part-1 during inference is not explicitly defined. To this end, we follow the Multi-head chen2023flexibility approach and choose the part with the most confident output (see below).
@@
Multi-head FL was proposed in chen2023flexibility and incorporated the multi-head approach into FL. Specifically, it considers a continual learning (CL) setup with incremental learning. Hence, a multi-head neural network architecture with task-specific layers is proposed to capture and preserve unique features for each task, enhancing prediction accuracy and model robustness. During the training phase, each client trains the whole model and the multi-heads locally. Then, at the end of each round, all layers are aggregated while preserving the multi-heads (i.e., the head $x$ of client $c$ is aggregated with the corresponding head from the other clients). As a result, the final model is comprised of two parts, the main body, shared for all tasks, and the different head layers. In order to make predictions during inference (when training has been completed), the following formula is used to make the selection between the head layers,
@@ C.1 Catastrophic Forgetting in Continual Learning
The goal of CL is to learn through a dynamic data distribution, and forgetting is measured by the metric of backward transfer, which estimates the impact of learning a task on the knowledge gained through the old tasks wang2024comprehensive. Mitigation techniques usually fall under 5 categories: regularization-based (discussed in Appendix C.2), replay-based, optimization-based, representation-based, and architecture-based. Works that provide insights on the phenomenon of forgetting through systematic analysis and propose mitigation techniques include ramasesh2020anatomy; ramasesh2021effect; kemker2018measuring. An aspect that is less studied in the literature, but relevant to our topic, is the impact of repetitions or cyclic order of the input data. In fact, repetitions can alleviate forgetting stojanov2019incremental; cossu2022class, and mitigation techniques based on multiple (sub-) networks per task/experience have proven effective against forgetting hemati2025continual. Theoretical results on cyclic learning for linear tasks are provided in evron2023continual; swartworth2023nearly. Another direction is to study CF with respect to the order of the input data in the data stream. Theoretical works focusing on linear models include lin2023theory. Finally, although not the focus of this manuscript, Curriculum Learning bengio2009curriculum studies the order of training samples with respect to their semantic relationship.
@@ C.2 Regularization in Continual Learning
In general, the concept of regularization involves adding a penalty term to the loss function, which restricts the weights from increasing (limiting the large changes in the weights). A common regularization technique in Machine Learning is L1 and L2 regularization, which is used to avoid model overfitting. Specifically, in the CL literature, the idea of regularization is extended as a way to balance old and new tasks by storing important model weights before altering them. One such example is EWC kirkpatrick2017overcoming, which calculates the importance of weights using the Fisher information matrix (FIM) and, according to that, adds a quadratic penalty in the loss function.
@@ D.1 Implementation Details
Repeated Experiments and Statistical Significance. All experiments are repeated at least $10$ times [^10] The only exception is when we study a single experiment, where we explicitly specify that it is a snapshot of a single experiment, as in Appendix G.2, for example. (changing seed). This is a common practice to avoid making assessments based on one experiment, see, e.g., li2022federated. As a result, throughout the paper, we report the median of the repeated experiments so that our results are robust against data outliers leys2013detecting. We note that variations in experimental results may occur due to weight initialization, stochastic training parameters, etc. Since historically, outliers were detected through mean-based metrics leys2013detecting, we also include the standard deviation. At the end of each round, we compute the accuracy of the global model (consisting of the aggregated part-1 and server’s part-2) and the PG, based on the per-label accuracies. Then, for each quantity, the median and standard deviation from all repeated experiments are computed.
@@ D.2 Cut layer selection
The cut layer is one of the key parameters in SL protocols since, as we show in our analysis, it affects the performance of the model when the training contains sequential updates between the clients (as in, e.g., SplitNN vepakomma2018split and SFL) and the impact of catastrophic forgetting. Moreover, the cut layer selection has an impact on the training time which, however, is not the focus of this work.
Considering the corresponding literature on how the cut layer selection is made, we notice that there are different approaches. Firstly, the authors in thapa2022splitfed selects arbitrarily relatively shallow cut layers, i.e., the second layer of LeNet (after the 2D MaxPool layer), the second layer of AlexNet krizhevsky2012imagenet (after the 2D MaxPool layer), the fourth layer of VGG16 simonyan2014very (after the 2D MaxPool layer), and the third layer (after the 2D BatchNormalization layer) of ResNet18 he2016deep. A similar approach is observed in liu2022wireless. For consistency throughout this manuscript, we consider the bottleneck search selection, which is common and widely used. This selection method was first introduced in kang2017neurosurgeon and has been applied in plenty of works of split computing, in general.
Finally, we note that the cut layer does not play a crucial role in SplitFedV1 in terms of performance. As we discussed in Appendix B, in SplitFedV1, all clients have a different version of the model parts that are trained in parallel (part-1 at the client and part-2 at the server). As a result, the performance of SplitFedV1 is the same as FL independently of the cut layer thapa2022splitfed.
@@ Appendix E Simulating Data Heterogeneity Patterns
Previous studies kairouz2021advances; zhu2021federated; solans2024non provide a taxonomy of the state-of-the-art strategies used to create non-IID data distributions for distributed learning research. One of the most prominent approaches among the main categories in the provided taxonomies corresponds to Label Distribution Skew, which implies that the label distribution $P_{i}(y)$ varies among clients while the conditional label-feature distribution $P_{i}(x|y)$ remains unchanged. Label distribution skewness seems to be the most challenging one li2022federated; liao2023mergesfl, especially when the clients have a restricted view of the training task. As discussed in previous sections of this manuscript, we conduct our experiments using Label Distribution Skew scenarios. In the upcoming subsections, we provide further details for each of the Label skew partitioning approaches that have been used.

### v2

@@ A.1 Algorithm of Hydra
Algorithm 1 presents in detail the workflow of Hydra. [^3] We note that part of this workflow is the same as (or aligns with) the workflow of SFL, e.g., how part-1 of the model is processed. For this reason, we often refer to it as SFL+Hydra. The algorithm describes the steps of the three main entities (i.e., Aggregator, Server, and Clients). In detail, the aggregator initiates the training by initializing the weights of the model, and sends the corresponding model parts to the server and the clients, i.e., weights $w_{p1}$ of part-1 are sent to clients, and weights $w_{p2_{a}},w_{p2_{b}}$ of part-2a and part-2b, are sent to the server. Then the server and the client can start training.
@@
At the beginning of the training, the group of clients corresponding to each head (note that the server maintains $G$ heads) is defined to group clients with similar data distributions. This grouping requires each client to send information on the distribution of labels in its local data before the training starts. The mapping (or grouping) of each client’s intermediate data to the different part-2b heads is done by using the $u$ set, which contains the mapping of clients into groups. This set is defined by Algorithm 2, following the assignment policy described in Appendix A.2.
Next, the training takes place in $R$ rounds. In every round, the server ensures that it receives the batch updates from all clients. Note that the server ‘pulls’ a new batch request (which contains the activation from a client) using the function proc_order(). This function guarantees that the preferred processing order is followed, e.g., cyclic order (or cyclic-and-reverse, as in Appendix F.2) or random order.
@@
Finally, at the end of each round, the aggregation takes place. In detail, the clients send the updated weights of part-1 to the aggregator, which will compute the weights for the next round, using an aggregation function, like FedAvg. We note that Hydra’s design provides the flexibility to apply any aggregation function, as it does not alter the processing workflow. Concurrently, the server aggregates locally all the different heads.
A.2 The Grouping Algorithm of Hydra
@@
In this section, we describe in detail how clients’ data are mapped to different versions of part-2b (heads). In its core, this mapping groups clients according to the similarity in their data. Therefore, we often use the terms “heads” and “groups” interchangeably. Let $g\in\mathcal{G}$ represent a group, where $\mathcal{G}$ is the set of all groups, and let $G=|\mathcal{G}|$ denote the total number of groups (each group corresponds to a part-2b head, as depicted in Figure 4). Hydra’s design naturally bounds $G$ by the number of different labels, i.e., $G\leq L$. This is in line with the multi-head architectures in CL, where, typically, the number of heads is equal to the number of tasks. In what follows, for simplicity, we describe the grouping algorithm of Hydra assuming that $G=L$. However, this description can be easily extended for the case where $G<L$, as we do in our ablation study in Section 5.
@@
The assignment of clients into groups considers two factors, (factor-i) the data distribution of the clients and, specifically, the representation of the different labels in the clients’ data, and (factor-ii) repartition of clients among the groups, i.e., ensuring that all heads receive a uniform number of updates (or, in other words, the groups have a balanced number of assigned clients). Regarding (factor-i), the goal is to group clients with similar data distributions. In this way, we avoid generating large gradients in the higher layers during iterations through clients. Recall that large gradients impose larger weight changes, which lead to overwriting previous updates, i.e., forgetting the previous knowledge. As shown in (Ramasesh et al., 2021a) (and verified in our observations), the higher layers are responsible for the large gradients that are back-propagated across the whole model. On the contrary, maintaining homogeneity within a group ensures a smooth transition between clients, minimizing significant weight changes and effectively addressing the primary challenge of CF.
Algorithm 2 Client assignment to groups (grouping algorithm)
Input: num. of groups $G$, set of clients $\mathcal{C}$
@@
Based on the discussion above, we propose the following approach for grouping the clients into groups and mapping them into different part-2b heads (as shown in Figure 4). Let $\rho_{cl}$ be the number of samples of label $l$ in client $c$, where $l\in\mathcal{L}$ and $c\in\mathcal{C}$, and $\rho_{c}=\{\rho_{cl}|l\in\mathcal{L}\},\forall c\in\mathcal{C}$ be the corresponding vector for each client $c$. Furthermore, we define the binary decision variables $\mathbf{u}=\{u_{cg}\mid\forall c\in\mathcal{C},\forall g\in\mathcal{G}\}$, where $u_{cg}=1$ when client $c$ is assigned to group $g$, and $u_{cg}=0$ otherwise. By design, each client must be assigned to exactly one group, i.e.,
@@ Appendix B Variants of Split (Federated) Learning
There is a plethora of proposed definitions for SL, e.g., SplitNN (Vepakomma et al., 2018) or parallel SL (Jeon and Kim, 2020), that mostly differ in the way the training at the server takes place and the presence of an aggregator. Among them, SplitFed (Thapa et al., 2022) (also called Split Federated Learning [^4] We clarify that, for simplicity, the abbreviation “SFL” always refers to SplitFedv2 throughout this manuscript, unless otherwise specified.) is an emerging method for training deep Neural Network models in a system comprised of multiple IoT and mobile devices. We introduce here the most widely used variants of SplitFed and discuss their differences. Figure 7 provides an overview of each variant of SplitFed, as well as the multi-head FL paradigm (Chen et al., 2023). We note that the training steps (forward/backward propagation) for the three variants of SplitFed (i.e., versions 1 to 3) are similar to the ones depicted in Figure 1. Moreover, in all three variants, part-1 is handled by the clients and part-2 is handled by the server.
@@
SplitFedV1 was the first SplitFed variant proposed in (Thapa et al., 2022), also often referred as Parallel SL (Tirana et al., 2024). As shown in Figure 7, there is a different version of the model parts for each client. This enables the distinct segments of the model to be trained entirely in parallel, replicating a paradigm similar to the one of FL. Additionally, the model parts are aggregated at the end of each round, and the global model is the result of concatenating the two model parts. The main drawback of this setup is the computing and memory demands on the server side, i.e., the server preserves the state of part-2 for every participating client. Moreover, the fact that the model convergence guarantees are the same as in FL, and not as in conventional SL, i.e., SplitNN (Vepakomma et al., 2018), can be considered another drawback. This is because often, as we have seen in the literature, SL paradigms can outperform FL (which is also shown in Figure 16 in the case of IID data distribution).
@@
SplitFedV2 (SFL) is the second variant of SplitFed introduced in (Thapa et al., 2022). This version aligns more with the original (sequential) Split Learning (SL) framework, i.e., SplitNN (Vepakomma et al., 2018), when compared to SplitFedV1. Unlike SplitFedV1, which operates with fully parallel model updates, SFL incorporates a shared model segment, specifically part-2, among the clients. Consequently, clients perform parallel updates on part-1, while part-2 is updated sequentially. The advantages of this approach include: (i) enhanced model accuracy relative to SplitFedV1 (FL), as shown in (Thapa et al., 2022) and (ii) reduced server resource requirements compared to SplitFedV1. However, as discussed in this paper, SFL is susceptible to catastrophic forgetting (CF), a phenomenon that is further pronounced under heterogeneous data distributions among the clients.
@@
SplitFedV3 was introduced  (Gawali et al., 2021) and (Madaan et al., 2022), and was specifically designed to address CF (Madaan et al., 2022) in Split Federated Learning (SFL). The training process closely mirrors that of SplitFedV1, where each client maintains its version of part-1 and part-2. However, at the end of each round, only part-2 is aggregated. Consequently, the final model features multiple input layers, corresponding to the layers that comprise part-1. Nevertheless, the process for selecting a part-1 during inference is not explicitly defined. To this end, we follow the Multi-head (Chen et al., 2023) approach and choose the part with the most confident output (see below).
@@
Multi-head FL was proposed in (Chen et al., 2023) and incorporated the multi-head approach into FL. Specifically, it considers a continual learning (CL) setup with incremental learning. Hence, a multi-head neural network architecture with task-specific layers is proposed to capture and preserve unique features for each task, enhancing prediction accuracy and model robustness. During the training phase, each client trains the whole model and the multi-heads locally. Then, at the end of each round, all layers are aggregated while preserving the multi-heads (i.e., the head $x$ of client $c$ is aggregated with the corresponding head from the other clients). As a result, the final model is comprised of two parts, the main body, shared for all tasks, and the different head layers. In order to make predictions during inference (when training has been completed), the following formula is used to make the selection between the head layers,
@@ C.1 Catastrophic Forgetting in Continual Learning
The goal of CL is to learn through a dynamic data distribution, and forgetting is measured by the metric of backward transfer, which estimates the impact of learning a task on the knowledge gained through the old tasks (Wang et al., 2024). Mitigation techniques usually fall under 5 categories: regularization-based (discussed in Appendix C.2), replay-based, optimization-based, representation-based, and architecture-based. Works that provide insights on the phenomenon of forgetting through systematic analysis and propose mitigation techniques include (Ramasesh et al., 2021a, b; Kemker et al., 2018). An aspect that is less studied in the literature, but relevant to our topic, is the impact of repetitions or cyclic order of the input data. In fact, repetitions can alleviate forgetting (Stojanov et al., 2019; Cossu et al., 2022), and mitigation techniques based on multiple (sub-) networks per task/experience have proven effective against forgetting (Hemati et al., 2025). Theoretical results on cyclic learning for linear tasks are provided in (Evron et al., 2023; Swartworth et al., 2023). Another direction is to study CF with respect to the order of the input data in the data stream. Theoretical works focusing on linear models include (Lin et al., 2023). Finally, although not the focus of this manuscript, Curriculum Learning (Bengio et al., 2009) studies the order of training samples with respect to their semantic relationship.
@@ C.2 Regularization in Continual Learning
In general, the concept of regularization involves adding a penalty term to the loss function, which restricts the weights from increasing (limiting the large changes in the weights). A common regularization technique in Machine Learning is L1 and L2 regularization, which is used to avoid model overfitting. Specifically, in the CL literature, the idea of regularization is extended as a way to balance old and new tasks by storing important model weights before altering them. One such example is EWC (Kirkpatrick et al., 2017), which calculates the importance of weights using the Fisher information matrix (FIM) and, according to that, adds a quadratic penalty in the loss function.
@@ D.1 Implementation Details
Repeated Experiments and Statistical Significance. All experiments are repeated at least $10$ times [^9] The only exception is when we study a single experiment, where we explicitly specify that it is a snapshot of a single experiment, as in Appendix G.2, for example. (changing seed). This is a common practice to avoid making assessments based on one experiment, see, e.g., (Li et al., 2022). We note that variations in experimental results may occur due to weight initialization, stochastic training parameters, etc. As a result, throughout the paper, we report the median of the repeated experiments so that our results are robust against outliers (Leys et al., 2013). Since historically, outliers were detected through mean-based metrics (Leys et al., 2013), we also include the standard deviation. Specifically, at the end of each round, we compute the accuracy of the global model (consisting of the aggregated part-1 and server’s part-2), the BW, and the PG, based on the per-label accuracies. Then, for each quantity, the median and standard deviation from all repeated experiments are computed.
@@ D.2 Cut Layer Selection
The cut layer is one of the key parameters in SL protocols since, as we show in our analysis, it affects the performance of the model when the training contains sequential updates between the clients (as in, e.g., SplitNN (Vepakomma et al., 2018) and SFL) and the impact of catastrophic forgetting. Moreover, the cut layer selection has an impact on the training time which, however, is not the focus of this work.
Considering the corresponding literature on how the cut layer selection is made, we notice that there are different approaches. Firstly, the authors in (Thapa et al., 2022) selects arbitrarily relatively shallow cut layers, i.e., the second layer of LeNet (after the 2D MaxPool layer), the second layer of AlexNet (Krizhevsky et al., 2012) (after the 2D MaxPool layer), the fourth layer of VGG16 (Simonyan and Zisserman, 2014) (after the 2D MaxPool layer), and the third layer (after the 2D BatchNormalization layer) of ResNet18 (He et al., 2016). A similar approach is observed in (Liu et al., 2022). For consistency throughout this manuscript, we consider the bottleneck search selection, which is common and widely used. This selection method was first introduced in (Kang et al., 2017) and has been applied in plenty of works of split computing, in general.
Finally, we note that the cut layer does not play a crucial role in SplitFedV1 in terms of performance. As we discussed in Appendix B, in SplitFedV1, all clients have a different version of the model parts that are trained in parallel (part-1 at the client and part-2 at the server). As a result, the performance of SplitFedV1 is the same as FL independently of the cut layer (Thapa et al., 2022).
@@ Appendix E Simulating Data Heterogeneity Patterns
Previous studies (Kairouz et al., 2021; Zhu et al., 2021; Jimenez G. et al., 2024) provide a taxonomy of the state-of-the-art strategies used to create non-IID data distributions for distributed learning research. One of the most prominent approaches among the main categories in the provided taxonomies corresponds to Label Distribution Skew, which implies that the label distribution $P_{i}(y)$ varies among clients while the conditional label-feature distribution $P_{i}(x|y)$ remains unchanged. Label distribution skewness seems to be the most challenging one (Li et al., 2022; Liao et al., 2024), especially when the clients have a restricted view of the training task. As discussed in previous sections of this manuscript, we conduct our experiments using Label Distribution Skew scenarios. In the upcoming subsections, we provide further details for each of the Label skew partitioning approaches that have been used.